OpenStudy (anonymous):

If 20 waves of water pass you in 10.0 s and the waves are 2.0 m long, how fast are they moving? A. 1.0 m/s B. 2.0 m/s C. 4.0 m/s D. 8.0 m/s ***Thank you!! :D

OpenStudy (michele_laino):

I think that we have to compute this quantity: \[\Large v = \frac{{20 \times 2}}{{10}} = ...\] since 1 wave passes at eac 0.5 seconds

OpenStudy (michele_laino):

each*

OpenStudy (anonymous):

okay! so 40/10 = 4 ? so the solution is choice C?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

yay!! thank you:)

OpenStudy (michele_laino):

thank you!! :)